# Solutions of the singular Yamabe problem near singular boundaries

### v1

@@ 2. Preliminaries
For some constants $A_{0},A_{1}>0$ and some function $\xi_{1}\in C^{\infty}(\Sigma)$ that satisfies (4.3) and (4.4) to be determined, we set
$w=u_{V}+\xi_{1}r^{-\frac{n}{2}+2}+A_{0}u_{V}r^{2}+A_{1}\phi_{1}r^{-\frac{n}{2}+3}.$
By (1.5) and $L_{1}\phi_{1}+\lambda_{1}\phi_{1}=0$, a direct computation yields
$\Delta w-\frac{1}{4}n(n-2)w^{\frac{n+2}{n-2}}$ $=\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}\Big(1+\frac{\xi_{1}}{\xi}r+A_{0}r^{2}+A_{1}\frac{\phi_{1}}{\xi}r^{2}\Big)^{\frac{n+2}{n-2}}$
$\leq\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}\Big[1+\frac{n+2}{n-2}\Big(\frac{\xi_{1}}{\xi}r+A_{0}r^{2}+A_{1}\frac{\phi_{1}}{\xi}r^{2}\Big)\Big]$
$=\Big(\frac{n(4-n)}{4}\xi_{1}+\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}\Big)r^{-\frac{n}{2}}$
$\qquad+A_{0}(4\xi-n\xi^{\frac{n+2}{n-2}})r^{-\frac{n}{2}+1}$
$\qquad+A_{1}\Big(\Big(-\frac{n}{2}+3\Big)\Big(\frac{n}{2}+1\Big)-\lambda_{1}\Big)\phi_{1}r^{-\frac{n}{2}+1}.$
@@
$a_{ij}=\delta_{ij}+a_{ij,k}y_{k}+O(r^{2}),\quad b_{i}=b_{i,0}+O(r).$
Hence, by (2.5), (2.6), (4.3),(4.4), and (A.12), we get
$\mathcal{L}w-\frac{n(n-2)}{4}w^{\frac{n+2}{n-2}}$
@@
Note that $F=-r^{\frac{n}{2}}(a_{ij,k}y_{k}(u_{V})_{ij}+b_{i,0}(u_{V})_{i})\in C^{\infty}(\Sigma)$ is a linear combination of $\nabla_{\theta}\xi$ and $\nabla^{2}_{\theta}\xi$ that satisfies (4.2). By applying Lemma 4.1 to such $F$, we obtain the desired $\xi_{1}$ satisfying (4.1). Therefore, we have
@@
$\leq\Big(\frac{n(4-n)}{4}\xi_{1}+\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}-F\Big)r^{-\frac{n}{2}}$
@@
Step 1. Set $A_{1}=kA_{0}$. First, we have $|O(1)|\leq C_{1}$ for $r\leq r_{0}$. Then, we choose $A_{0}>\max\{2C_{1}/n,1\}$. We claim that by choosing $k>0$ large enough, we have
@@
Recall (1.14). Note that $\mu_{1}>2$ implies that $(-n/2+3)(n/2+1)-\lambda_{1}<0$. Hence, the left-hand side of (4.12) is controlled by
$C_{1}\xi^{\frac{n+2}{n-2}}+A_{0}(4\xi-n\xi^{\frac{n+2}{n-2}})\leq A_{0}\Big(\frac{n}{2}\xi^{\frac{n+2}{n-2}}+4\xi-n\xi^{\frac{n+2}{n-2}}\Big)\leq-\frac{n}{4}A_{0}\xi^{\frac{n+2}{n-2}}\quad\text{on }\Sigma_{\delta}.$
@@
for some positive constants $c_{\delta}$ and $C_{\delta}$. Hence, the left-hand side of (4.12) is controlled by
@@
on $\Sigma\backslash\Sigma_{\delta}$, by taking $k$ such that
(4.13) $k>\Big(\frac{n}{2}C^{\frac{n+2}{n-2}}_{\delta}+4C_{\delta}\Big)\Big/\Big(\Big(\frac{n}{2}-3\Big)\Big(\frac{n}{2}+1\Big)+\lambda_{1}\Big)c_{\delta}.$
In summary, by choosing $A_{0}>\max\{2C_{1}/n,1\}$ and $k$ satisfying (4.13), we obtain (4.12).
@@
(4.14) $(A_{0}+A_{1}+1)O(r^{-\frac{n}{2}+2})\leq(k+2)A_{0}rO(r^{-\frac{n}{2}+1})\leq\frac{n}{8}A_{0}r^{-\frac{n}{2}+1}\quad\text{for any }r\leq r^{\prime}_{0}.$
@@
In summary, set $\bar{r}_{1}=\min\{r_{0},r^{\prime}_{0},1/(2C_{2})\}$ and take $A_{0}>\max\{2C_{1}/n,1,(B+C_{2})/\bar{r}_{1}\}$ and $A_{1}=kA_{0}$ with (4.13). By the construction of $\xi_{1}$, (4.12), and (4.14), we have for any $r\leq\bar{r}_{1}$,
@@
In addition, by (4.15), we have $w=+\infty$ on $\partial V\cap\{r\leq\bar{r}_{1}\}$. On the other hand, by (4.11),(4.15), and the choose of $A_{0}$, we get, on $r=\bar{r}_{1}$,
@@
For some constants $A_{0},A_{1}>0$ and some function $\xi_{1}\in C^{\infty}(\Sigma)$ that satisfies (4.3) and (4.4) to be determined, we set
$w=u_{V}+\xi_{1}r^{-\frac{n}{2}+2}+A_{0}u_{V}r^{2}-A_{1}\phi_{1}r^{-\frac{n}{2}+3}\log r.$
@@
$=\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}\Big(1+\frac{\xi_{1}}{\xi}r+A_{0}r^{2}-A_{1}\frac{\phi_{1}}{\xi}r^{2}\log r\Big)^{\frac{n+2}{n-2}}$
$\leq\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}\Big[1+\frac{n+2}{n-2}\Big(\frac{\xi_{1}}{\xi}r+A_{0}r^{2}-A_{1}\frac{\phi_{1}}{\xi}r^{2}\log r\Big)\Big]$
$=\Big(\frac{n(4-n)}{4}\xi_{1}+\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}\Big)r^{-\frac{n}{2}}$
@@
$\qquad+A_{0}(4\xi-n\xi^{\frac{n+2}{n-2}})r^{-\frac{n}{2}+1}-4A_{1}\phi_{1}r^{-\frac{n}{2}+1}.$
@@
$\leq\Big(\frac{n(4-n)}{4}\xi_{1}+\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}-F\Big)r^{-\frac{n}{2}}$
@@
$\qquad+(A_{0}+A_{1}+1)\xi^{\frac{n+2}{n-2}}O(r^{-\frac{n}{2}+2}\log r).$
Proceeding as in Case 1, we can choose $\xi_{1}$, $A_{0}$, $A_{1}$, and $r_{1}$ such that
@@
For some constants $A_{0},A_{1}>0$ and some function $\xi_{1}\in C^{\infty}(\Sigma)$ that satisfies (4.3) and (4.4) to be determined, we set
$w=u_{V}+\xi_{1}r^{-\frac{n}{2}+2}+A_{0}u_{V}r^{2}+A_{1}\phi_{1}(r^{\mu_{1}-\frac{n-2}{2}}-r^{-\frac{n}{2}+3}).$
By $L_{1}\phi_{1}+\lambda_{1}\phi_{1}=0$ and (1.14), we have
@@
By a computation similar to Case 1, we have
@@
$=\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}\Big(1+\frac{\xi_{1}}{\xi}r+A_{0}r^{2}+A_{1}\frac{\phi_{1}}{\xi}(r^{\mu_{1}}-r^{2})\Big)^{\frac{n+2}{n-2}}$
$\leq\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}\Big[1+\frac{n+2}{n-2}\Big(\frac{\xi_{1}}{\xi}r+A_{0}r^{2}+A_{1}\frac{\phi_{1}}{\xi}(r^{\mu_{1}}-r^{2})\Big)\Big]$
$=\Big(\frac{n(4-n)}{4}\xi_{1}+\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}\Big)r^{-\frac{n}{2}}+A_{0}(4\xi-n\xi^{\frac{n+2}{n-2}})r^{-\frac{n}{2}+1}$
$\qquad-A_{1}\Big(\Big(-\frac{n}{2}+3\Big)\Big(\frac{n}{2}+1\Big)-\lambda_{1}\Big)\phi_{1}r^{-\frac{n}{2}+1}.$
Note that $\mu_{1}<2$ implies that $(-n/2+3)(n/2+1)-\lambda_{1}>0$. By (1.14), Lemma A.2(1), and Lemma A.2(3), we have $\mu_{1}>1$ for any $n\geq 3$. Hence, we have
@@
$\leq\Big(\frac{n(4-n)}{4}\xi_{1}+\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}-F\Big)r^{-\frac{n}{2}}$
@@
$\qquad+(A_{0}+A_{1}+1)\xi^{\frac{n+2}{n-2}}O(r^{-\frac{n}{2}+\mu_{1}}).$
Proceeding as in Case 1, we can choose $\xi_{1}$, $A_{0}$, $A_{1}$, and $r_{1}$ such that

### v2

@@ 2. Preliminaries
Let $A_{0},A_{1}>0$ be constants and let $\xi_{1}\in C^{\infty}(\Sigma)$ be a function to be determined satisfying estimates of the form (4.4) and (4.6), i.e., for some $\beta\in(0,1)$,
(4.14) $|\xi_{1}|+d_{\Sigma}|\nabla_{\theta}\xi_{1}|+d_{\Sigma}^{2}|\nabla^{2}_{\theta}\xi_{1}|+d_{\Sigma}^{2+\beta}[\nabla^{2}_{\theta}\xi_{1}]_{C^{\beta}(B_{d_{\Sigma}(\cdot)/2}(\cdot))}\leq C_{0}\xi,$
for some constants $C_{0}>0$ depends only on $n$, $\beta$, $g$, $V$, and the derivatives of $T$ at 0. We set
$w=u_{V}+\xi_{1}r^{-\frac{n}{2}+2}+A_{0}u_{V}r^{2}+A_{1}\phi_{1}r^{-\frac{n}{2}+3}=u_{V}(1+R_{1}),$
where
$R_{1}:=\frac{\xi_{1}}{\xi}r+A_{0}r^{2}+A_{1}\frac{\phi_{1}}{\xi}r^{2}.$
We will choose $A_{0}$, $A_{1}$, $\xi_{1}$, and $\bar{r}_{1}>0$ small enough so that
(4.15) $C_{0}\bar{r}_{1}+A_{0}\bar{r}^{2}_{1}+A_{1}C^{\prime}\bar{r}^{2}_{1}\leq\frac{1}{2},$
hence we have, for any $r\leq\bar{r}_{1}$,
(4.16) $\left|(1+R_{1})^{\frac{n+2}{n-2}}-1-\frac{n+2}{n-2}R_{1}\right|\leq C(n)R^{2}_{1},$
where $C(n)>0$ depends only on $n$.
We suppose (4.15) holds, then by (1.5) and $L_{1}\phi_{1}+\lambda_{1}\phi_{1}=0$, a direct computation yields
(4.17) $\Delta w-\frac{1}{4}n(n-2)w^{\frac{n+2}{n-2}}$
$=\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}(1+R_{1})^{\frac{n+2}{n-2}}$
$\leq\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}\Big[1+\frac{n+2}{n-2}R_{1}-C(n)R^{2}_{1}\Big]$
$=\Big(\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}+\frac{n(4-n)}{4}\xi_{1}\Big)r^{-\frac{n}{2}}+A_{0}(4\xi-n\xi^{\frac{n+2}{n-2}})r^{-\frac{n}{2}+1}$
$\qquad+A_{1}\Big(\Big(-\frac{n}{2}+3\Big)\Big(\frac{n}{2}+1\Big)-\lambda_{1}\Big)\phi_{1}r^{-\frac{n}{2}+1}+\frac{1}{4}n(n-2)C(n)u_{V}^{\frac{n+2}{n-2}}R^{2}_{1}.$
@@
$a_{ij}=\delta_{ij}+a_{ij,k}y_{k}+O(r^{2}),\mbox{ and }\ b_{i}=b_{i,0}+O(r).$
Hence, by (2.5), (2.6), (4.14), and (A.12), we get
(4.18) $\mathcal{L}w-\frac{n(n-2)}{4}w^{\frac{n+2}{n-2}}$
@@
Note that $F_{1}:=-r^{\frac{n}{2}}(a_{ij,k}y_{k}(u_{V})_{ij}+b_{i,0}(u_{V})_{i})\in C^{\infty}(\Sigma)$ is a linear combination of $\nabla_{\theta}\xi$ and $\nabla^{2}_{\theta}\xi$. It follows from (2.6) that $F_{1}$ satisfies (4.3) and (4.5). By Remark 4.2, we can apply Lemma 4.1(i) with $F=F_{1}$ and $k=1$ to obtain the desired $\xi_{1}$ satisfying (4.1) with $\sigma=0$, and (4.14). Therefore, by substituting (4.17) into (4.18), we have
@@
$\leq\Big(\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}+\frac{n(4-n)}{4}\xi_{1}-F_{1}\Big)r^{-\frac{n}{2}}$
@@
where the first term on the right hand side vanishes.
Step 1. Set $A_{1}=kA_{0}$. First, we have $|O(1)|\leq C_{1}$ for $r\leq r_{0}$. Then, we choose $A_{0}\geq\max\{2C_{1}/n,1\}$. We claim that by choosing $k>0$ large enough, we have
@@
Recall (1.14), and note that $\mu_{1}>2$ implies $(-n/2+3)(n/2+1)-\lambda_{1}<0$. Hence, on $\Sigma_{\delta}$ the left-hand side of (4.19) is controlled by
$C_{1}\xi^{\frac{n+2}{n-2}}+A_{0}(4\xi-n\xi^{\frac{n+2}{n-2}})\leq A_{0}\Big(\frac{n}{2}\xi^{\frac{n+2}{n-2}}+4\xi-n\xi^{\frac{n+2}{n-2}}\Big)\leq-\frac{n}{4}A_{0}\xi^{\frac{n+2}{n-2}}.$
@@
for some positive constants $c_{\delta}$ and $C_{\delta}$. Hence, on $\Sigma\backslash\Sigma_{\delta}$ the left-hand side of (4.19) is controlled by
@@
where we take
(4.20) $k=\Big(\frac{n}{2}C^{\frac{n+2}{n-2}}_{\delta}+4C_{\delta}\Big)\Big/\Big(\Big(\frac{n}{2}-3\Big)\Big(\frac{n}{2}+1\Big)+\lambda_{1}\Big)c_{\delta}.$
In summary, by choosing $A_{0}\geq\max\{2C_{1}/n,1\}$ and $k$ as (4.20), we obtain (4.19).
@@
(4.21) $(A_{0}+A_{1}+1)O(r^{-\frac{n}{2}+2})$ $\leq(k+2)A_{0}r|O(r^{-\frac{n}{2}+1})|$
$\leq\frac{n}{8}A_{0}r^{-\frac{n}{2}+1}\quad\text{for any }r\leq r^{\prime}_{0}.$
@@
Recall $A_{1}=kA_{0}$ with $k$ as (4.20), now we set
$\bar{r}_{1}=\min\left\{1,r_{0},r^{\prime}_{0},\frac{1}{2C_{2}},\frac{1}{100C_{0}},\frac{1}{4(1+B+C_{2}+2C_{1}/n)(1+kC^{\prime})}\right\},$
and take
$A_{0}=\max\left\{\frac{2C_{1}}{n},1,\frac{B+C_{2}}{\bar{r}_{1}}\right\}.$
Hence, we have (4.15). By the construction of $\xi_{1}$, (4.19), and (4.21), we have for any $r\leq\bar{r}_{1}$,
@@
In addition, by (4.22), we have $w=+\infty$ on $\partial V\cap\{r\leq\bar{r}_{1}\}$. On the other hand, by (4.13), (4.22), and the choice of $A_{0}$, we get, on $\{r=\bar{r}_{1}\}$,
@@
For some constants $A_{0},A_{1}>0$ and some function $\xi_{1}\in C^{\infty}(\Sigma)$ that satisfies (4.14) to be determined, we set
$w=u_{V}+\xi_{1}r^{-\frac{n}{2}+2}+A_{0}u_{V}r^{2}-A_{1}\phi_{1}r^{-\frac{n}{2}+3}\log r=u_{V}(1+R_{1}),$
where
$R_{1}:=\frac{\xi_{1}}{\xi}r+A_{0}r^{2}-A_{1}\frac{\phi_{1}}{\xi}r^{2}\log r.$
We will choose $A_{0}$, $A_{1}$, $\xi_{1}$, and $\bar{r}_{1}\in(0,1/2)$ small enough so that
$C_{0}\bar{r}_{1}+A_{0}\bar{r}^{2}_{1}-A_{1}C^{\prime}\bar{r}^{2}_{1}\log\bar{r}_{1}\leq\frac{1}{2},$
then, for any $r\leq\bar{r}_{1}$, (4.16) holds for such $R_{1}$.
@@
$=\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}(1+R_{1})^{\frac{n+2}{n-2}}$
$\leq\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}\Big[1+\frac{n+2}{n-2}R_{1}-C(n)R^{2}_{1}\Big]$
$=\Big(\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}+\frac{n(4-n)}{4}\xi_{1}\Big)r^{-\frac{n}{2}}$
@@
$\qquad+A_{0}(4\xi-n\xi^{\frac{n+2}{n-2}})r^{-\frac{n}{2}+1}-4A_{1}\phi_{1}r^{-\frac{n}{2}+1}+\frac{1}{4}n(n-2)C(n)u_{V}^{\frac{n+2}{n-2}}R^{2}_{1}.$
@@
$\leq\Big(\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}+\frac{n(4-n)}{4}\xi_{1}-F_{1}\Big)r^{-\frac{n}{2}}$
@@
$\qquad+(A_{0}+A_{1}+1)\xi^{\frac{n+2}{n-2}}O(r^{-\frac{n}{2}+2}\log r),$
where $F_{1}$ is the same as that in Case 1. Proceeding as in Case 1, we can choose $\xi_{1}$, $A_{0}$, $A_{1}$, and $r_{1}$ such that
@@
for any $x\in\Omega\cap B_{r_{1}}(0)$, by taking $r_{1}$ smaller. The above two inequalities imply the second part of (1.13).
Next, we improve this estimate to (1.15) if $T\in C^{3,\gamma}$, for some $\gamma\in(0,1)$. Note in this case, $a_{ij}\in C^{2,\gamma},b_{i}\in C^{1,\gamma}$, and we know
$a_{ij}=\delta_{ij}+a_{ij,k}y_{k}+a_{ij,kl}y_{k}y_{l}+O(r^{2+\gamma}),\mbox{ and }b_{i}=b_{i,0}+b_{i,k}y_{k}+O(r^{1+\gamma}).$
Let $A_{0},A_{1}>0$ be constants and let $\xi_{1},\xi_{2},\xi_{3}\in C^{\infty}(\Sigma)$ be functions to be determined that satisfy estimates of the form (4.4) and (4.6), i.e., for some $\beta\in(0,1)$,
(4.23) $|\xi_{i}|+d_{\Sigma}|\nabla_{\theta}\xi_{i}|+d_{\Sigma}^{2}|\nabla^{2}_{\theta}\xi_{i}|+d_{\Sigma}^{2+\beta}[\nabla^{2}_{\theta}\xi_{i}]_{C^{\beta}(B_{d_{\Sigma}(\cdot)/2}(\cdot))}\leq C_{0}\xi,\quad i=1,2,3,$
for some constants $C_{0}>0$ depends only on $n$, $\beta$, $g$, $V$, and the derivatives of $T$ at 0. We set
$w$ $=u_{V}+\xi_{1}r^{-\frac{n}{2}+2}+\xi_{2}r^{-\frac{n}{2}+3}\log r+\xi_{3}r^{-\frac{n}{2}+3}+A_{0}u_{V}r^{2+\gamma}+A_{1}\phi_{1}r^{-\frac{n}{2}+1}(r^{2}-r^{2+\gamma})$
$=u_{V}(1+R_{2}),$
where
$R_{2}:=\frac{\xi_{1}}{\xi}r+\frac{\xi_{2}}{\xi}r^{2}\log r+\frac{\xi_{3}}{\xi}r^{2}+A_{0}r^{2+\gamma}+A_{1}\frac{\phi_{1}}{\xi}(r^{2}-r^{2+\gamma}).$
We will choose $A_{0}$, $A_{1}$, $\xi_{i}(i=1,2,3)$, and $\bar{r}_{1}\in(0,1/2)$ small enough so that
$C_{0}(\bar{r}_{1}-\bar{r}_{1}^{2}\log\bar{r}_{1}+\bar{r}^{2}_{1})+A_{0}\bar{r}^{2+\gamma}_{1}+A_{1}C^{\prime}\bar{r}_{1}^{2}\leq\frac{1}{2},$
hence we have, for any $r\leq\bar{r}_{1}$,
$\left|(1+R_{2})^{\frac{n+2}{n-2}}-1-\frac{n+2}{n-2}R_{2}-\frac{2(n+2)}{(n-2)^{2}}R^{2}_{2}\right|\leq C(n)|R_{2}|^{3},$
where $C(n)>0$ depends only on $n$. Note that $\mu_{1}=2$ implies that $(-n/2+3)(n/2+1)-\lambda_{1}=0,$ then by $L_{1}\phi_{1}+\lambda_{1}\phi_{1}=0$, we have
$\Delta(\phi_{1}r^{-\frac{n}{2}+3})-\frac{n(n+2)}{4\rho^{2}}r^{-\frac{n}{2}+1}\phi_{1}=0.$
A computation similar to Case 1 yields
$\Delta w-\frac{1}{4}n(n-2)w^{\frac{n+2}{n-2}}$
$=\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}(1+R_{2})^{\frac{n+2}{n-2}}$
$\leq\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}\Big[1+\frac{n+2}{n-2}R_{2}+\frac{2(n+2)}{(n-2)^{2}}R^{2}_{2}-C(n)|R_{2}|^{3}\Big]$
$=\Big(\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}+\frac{n(4-n)}{4}\xi_{1}\Big)r^{-\frac{n}{2}}$
$\qquad+\Big(\Delta_{\theta}\xi_{2}-\frac{n(n+2)}{4\rho^{2}}\xi_{2}+\left(-\frac{n}{2}+3\right)\left(\frac{n}{2}+1\right)\xi_{2}\Big)r^{-\frac{n}{2}+1}\log r$
$\qquad+\Big(\Delta_{\theta}\xi_{3}-\frac{n(n+2)}{4\rho^{2}}\xi_{3}+\left(-\frac{n}{2}+3\right)\left(\frac{n}{2}+1\right)\xi_{3}+4\xi_{2}-\frac{n(n+2)}{2(n-2)}\xi^{\frac{n+2}{n-2}}\Big(\frac{\xi_{1}}{\xi}\Big)^{2}\Big)r^{-\frac{n}{2}+1}$
$\qquad+A_{0}\Big((2+\gamma)^{2}\xi-n\xi^{\frac{n+2}{n-2}}\Big)r^{-\frac{n}{2}+1+\gamma}-A_{1}\Big((2+\gamma)^{2}-2^{2}\Big)\phi_{1}r^{-\frac{n}{2}+1+\gamma}$
$\qquad+(A_{0}+A_{1}+1)\xi^{\frac{n+2}{n-2}}O(r^{-\frac{n}{2}+2}).$
Hence, we have
$\mathcal{L}w-\frac{n(n-2)}{4}w^{\frac{n+2}{n-2}}$
$=\Big(\Delta w-\frac{n(n-2)}{4}w^{\frac{n+2}{n-2}}\Big)+a_{ij,k}y_{k}w_{ij}+a_{ij,kl}y_{k}y_{l}w_{ij}+O(r^{2+\gamma})w_{ij}$
$\qquad+b_{i,0}w_{i}+b_{i,k}y_{k}w_{i}+O(r^{1+\gamma})w_{i}+cw$
$\qquad+a_{ij,kl}y_{k}y_{l}(u_{V})_{ij}+b_{i,k}y_{k}(u_{V})_{i}+cu_{V}+a_{ij,k}y_{k}(\xi_{1}r^{-\frac{n}{2}+2})_{ij}+b_{i,0}(\xi_{1}r^{-\frac{n}{2}+2})_{i}$
$\qquad+O(r^{2+\gamma})(u_{V})_{ij}+O(r^{1+\gamma})(u_{V})_{i}+\xi^{\frac{n+2}{n-2}}O(r^{-\frac{n}{2}+2}\log r)$
$\qquad+(A_{0}+A_{1}+1)\xi^{\frac{n+2}{n-2}}O(r^{-\frac{n}{2}+2})$
$\leq\Big(\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}+\frac{n(4-n)}{4}\xi_{1}-F_{1}\Big)r^{-\frac{n}{2}}$
$\qquad+\Big(\Delta_{\theta}\xi_{2}-\frac{n(n+2)}{4\rho^{2}}\xi_{2}+\left(-\frac{n}{2}+3\right)\left(\frac{n}{2}+1\right)\xi_{2}\Big)r^{-\frac{n}{2}+1}\log r$
$\qquad+\Big(\Delta_{\theta}\xi_{3}-\frac{n(n+2)}{4\rho^{2}}\xi_{3}+\left(-\frac{n}{2}+3\right)\left(\frac{n}{2}+1\right)\xi_{3}+4\xi_{2}-F_{2}\Big)r^{-\frac{n}{2}+1}$
$\qquad+\Big(\xi^{\frac{n+2}{n-2}}O(1)+A_{0}((2+\gamma)^{2}\xi-n\xi^{\frac{n+2}{n-2}})\Big)r^{-\frac{n}{2}+1+\gamma}-A_{1}\Big((2+\gamma)^{2}-2^{2}\Big)\phi_{1}r^{-\frac{n}{2}+1+\gamma}$
$\qquad+(A_{0}+A_{1}+1)\xi^{\frac{n+2}{n-2}}O(r^{-\frac{n}{2}+2}),$
where
$F_{1}$ $=-r^{\frac{n}{2}}\Big(a_{ij,k}y_{k}(u_{V})_{ij}+b_{i,0}(u_{V})_{i}\Big),$
$F_{2}$ $=-r^{\frac{n}{2}-1}\Big\{a_{ij,kl}y_{k}y_{l}(u_{V})_{ij}+b_{i,k}y_{k}(u_{V})_{i}+cu_{V}$
$\qquad+a_{ij,k}y_{k}(\xi_{1}r^{-\frac{n}{2}+2})_{ij}+b_{i,0}(\xi_{1}r^{-\frac{n}{2}+2})_{i}\Big\}+{\frac{n(n+2)}{2(n-2)}\xi^{\frac{n+2}{n-2}}\Big(\frac{\xi_{1}}{\xi}\Big)^{2}}.$
First, by the same reasoning as in Case 1, we can find a $\xi_{1}\in C^{\infty}(\Sigma)$ such that
$\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}+\frac{n(4-n)}{4}\xi_{1}-F_{1}$ $=0.$
Then we choose $\xi_{2}=-\frac{\sigma}{4}\phi_{1}$, where $\sigma$ comes from Lemma 4.1(ii) with $F=F_{2}$ and $k=2$ in (4.1). Finally, by Lemma 4.1(ii) with $F=F_{2}$ and $k=2$, we can find $\xi_{3}\in C^{\infty}(\Sigma)$ such that
$\Delta_{\theta}\xi_{3}-\frac{n(n+2)}{4\rho^{2}}\xi_{3}+\left(-\frac{n}{2}+3\right)\left(\frac{n}{2}+1\right)\xi_{3}+4\xi_{2}-F_{2}$ $=0.$
By Lemma 4.1 and Lemma A.2(4), we know that $\xi_{i}$ satisfies (4.23), $i=1,2,3$. Proceeding as in Case 1, we can choose $A_{0}$, $A_{1}$, and $r_{1}$ such that
$v\leq w\leq u_{V}+c_{1}u_{V}r+c_{2}u_{V}r^{2}\log r+Cu_{V}d^{2}_{g}(x,0),$
for any $x\in\Omega\cap B_{r_{1}}(0)$, where $c_{1}=\xi_{1}/\xi$ and $c_{2}=\xi_{2}/\xi$ are both in $C^{\infty}(\Sigma)\cap L^{\infty}(\Sigma)$. Set $\bar{w}=u_{V}+\xi_{1}r^{-\frac{n}{2}+2}+\xi_{2}r^{-\frac{n}{2}+3}\log r+\xi_{3}r^{-\frac{n}{2}+3}-A_{0}u_{V}r^{2+\gamma}-A_{1}\phi_{1}r^{-\frac{n}{2}+1}(r^{2}-r^{2+\gamma})$, a similar discussion yields
$v\geq\bar{w}\geq u_{V}+c_{1}u_{V}r+c_{2}u_{V}r^{2}\log r-Cu_{V}d^{2}_{g}(x,0),$
@@
For some constants $A_{0},A_{1}>0$ and some function $\xi_{1}\in C^{\infty}(\Sigma)$ satisfying (4.14) to be determined, we set
$w=u_{V}+\xi_{1}r^{-\frac{n}{2}+2}+A_{0}u_{V}r^{2}+A_{1}\phi_{1}(r^{\mu_{1}-\frac{n-2}{2}}-r^{-\frac{n}{2}+3})=u_{V}(1+R_{1}),$
where
$R_{1}:=\frac{\xi_{1}}{\xi}r+A_{0}r^{2}+A_{1}\frac{\phi_{1}}{\xi}(r^{\mu_{1}}-r^{2}).$
We will choose $A_{0}$, $A_{1}$, $\xi_{1}$, and $\bar{r}_{1}>0$ small enough so that
$C_{0}\bar{r}_{1}+A_{0}\bar{r}^{2}_{1}+A_{1}C^{\prime}\bar{r}^{\mu_{1}}_{1}\leq\frac{1}{2},$
then, for any $r\leq\bar{r}_{1}$, (4.16) holds for such $R_{1}$. By $L_{1}\phi_{1}+\lambda_{1}\phi_{1}=0$ and (1.14), we have
@@
By a computation similar to Case 1, we get
@@
$=\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}(1+R_{1})^{\frac{n+2}{n-2}}$
$\leq\Delta w-\frac{1}{4}n(n-2)u_{V}^{\frac{n+2}{n-2}}\Big[1+\frac{n+2}{n-2}R_{1}-C(n)R^{2}_{1}\Big]$
$=\Big(\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}+\frac{n(4-n)}{4}\xi_{1}\Big)r^{-\frac{n}{2}}+A_{0}(4\xi-n\xi^{\frac{n+2}{n-2}})r^{-\frac{n}{2}+1}$
$\qquad-A_{1}\Big(\Big(-\frac{n}{2}+3\Big)\Big(\frac{n}{2}+1\Big)-\lambda_{1}\Big)\phi_{1}r^{-\frac{n}{2}+1}+\frac{1}{4}n(n-2)C(n)u_{V}^{\frac{n+2}{n-2}}R^{2}_{1}.$
Note that $\mu_{1}<2$ implies that $(-n/2+3)(n/2+1)-\lambda_{1}>0$. By (1.14), Lemma A.2(1), and Lemma A.2(3), we have $\mu_{1}>1$ for any $n\geq 3$. Hence, we obtain
@@
$\leq\Big(\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4\rho^{2}}\xi_{1}+\frac{n(4-n)}{4}\xi_{1}-F_{1}\Big)r^{-\frac{n}{2}}$
@@
$\qquad+(A_{0}+A_{1}+1)\xi^{\frac{n+2}{n-2}}O(r^{-\frac{n}{2}+\mu_{1}}),$
where $F_{1}$ is the same as in Case 1. Proceeding as in Case 1, we can choose $\xi_{1}$, $A_{0}$, $A_{1}$, and $r_{1}$ such that